A 306 inch board is cut into 2 pieces. One piece is five times the length of the other. Find the length of the two pieces.

The short piece is ___________
The long piece is ____________

Answer by mangopeeler07(462)

x=short piece
5x=long piece

5x+x=306------------------because the pieces together are 306 inches long

6x=306--------------------because we had to combine like terms

x=51----------------------because we had to divide both sides by 6 to isolate x

short piece=51 inches-----because short piece=x and x=51
long piece= 255 inches-----because long piece=5x and 5x=5(51)

5x+x=306.......255+51=306-----------------because we had to check our answer

Answer:
short piece=51 inches
long piece= 255 inches